OpenStudy (anonymous):

If P(x)=(x-a)Q(x), where Q(x) is a polynomial, explain why p(a)=0???

OpenStudy (lgbasallote):

because (x-a) is a factor...and according to remainder theorem if x-a is a factor then p(a) = 0

OpenStudy (anonymous):

or simply replace \(x\) by \(a\) and see what you get

OpenStudy (lgbasallote):

and also because if you sub a into x you get a-a = 0

OpenStudy (anonymous):

oh so because (a-a)=0, 0xQ(a)=0 also.

OpenStudy (lgbasallote):

yup